# Spin systems as quantum simulators of quantum field theories in curved spacetimes

### v1

@@ III.1 Spin models and their continuum limits
where $k$ is a typical wave number333 The long-wavelength limit of the spin system does not imply the low-energy limit in general. For example, in the XY-model, the gap closes twice in the momentum space and the Dirac fermion appears instead of the Majorana fermion in the low-energy limit. In this paper, we will focus on the long-wavelength limit and consider the Majorana field. . Since the physical system size is given by $\ell$, the typical wave number should satisfy $k\gtrsim 1/\ell$. Similarly, we can express all terms in Eq. (21) using $\Psi(x_{j})$. We will assume that $J^{ab}_{j}(t)=\mathcal{O}(\varepsilon^{-1})$, $h_{j}(t)=\mathcal{O}(\varepsilon^{-1})$ and $J^{+-}_{j}(t)+J^{-+}_{j}(t)-2h_{j}(t)=\mathcal{O}(\varepsilon^{0})$. In terms of their limiting behaviors, we define the real functions $v(t,x)$, $w(t,x)$, $p(t,x)$, $q(t,x)$ and $r(t,x)$ as follows:
@@ IV.3 Exactly solvable example
This satisfies $a(\eta)\to a_{1}$ ($\eta\to-\infty$) and $a(\eta)\to a_{2}$ ($\eta\to\infty$). In this case, we have analytical solutions of mode functions both in continuum and discrete theories. The detailed calculation is summarized in appendix B. Once we have two sets of mode functions, we can compute the number of produced particles for a given wave number using Eqs. (54) and (65). As the quantum state $|\Omega\rangle$, we take the ground state at the sufficiently early time. Figure 1 shows the number of particles produced by the cosmic expansion with $m=a_{1}=1$, $a_{2}=2$ and $\ell/2\pi=5$. The upper and lower panels are for $\Delta\eta=1$ and $\Delta\eta=0.2$. Both results for the continuum and discrete theories are shown together. For the discrete theory, we take $k=\kappa/\varepsilon$ as the horizontal axis for the visibility of the continuum limit $L\to\infty$. The number of sites is varied as $L=64,128,256,512$. Since we consider that $\ell$ is fixed, the lattice spacing $\varepsilon=\ell/L$ depends on $L$ while the wave numbers $k$ are defined at the same discrete points for any $L$. In Fig. 1, we plot $n_{k}$ as a function of the wave number at the discrete points represented by dots, together with a continuum variable for visibility. We find that the spectrum tends to converge to that of the continuum theory as $L$ increases. Indeed, the difference between the number of particles produced in discrete and continuum theories, $|n_{k}(L)-n_{k}^{\text{QFT}}|$, depends on $O(1/L)$, as shown in Fig. 2. Especially, for $L\gtrsim 100$, the spectrum qualitatively agrees with that of the continuum theory. This result demonstrates that the spin model can work as a simulator of QFT in the curved spacetime. As the timescale of the cosmic expansion $\Delta\eta$ becomes smaller, the more particles are produced. As the timescale of the cosmic expansion $\Delta\eta$ decreases, more particles are produced. Also, typical wave number of produced particles increases as $\Delta\eta$ decreases. Now, a physical length scale of the systems is given by $\ell$. A typical Hubble scale is $H_{\eta}\equiv a^{-2}da/d\eta\approx 1/a\Delta\eta$ during the cosmic expansion. Therefore, the present phenomena should be governed by the relation between the values of two dimensionless parameters, the mass scale $m\ell$ and the Hubble scale $H_{\eta}\ell$.
@@ V.3 Confirmation of the Unruh effect through the spin system
where we have used $I(r^{\prime})=(\pi/2)(1+r^{\prime}{}^{2}/4+\cdots)=(\pi/2)(1+\varepsilon m/2+\cdots)$. These are the metric functions of the Rindler spacetime with $\kappa=2\pi T_{U}$, as shown in (71). Thus, the entanglement Hamiltonian (69) is directly related to the Rindler Hamiltonian $H_{R}$ in the continuum limit and exhibits the Unruh effect.
@@ VI Summary and discussion
In this work, we established a complete dictionary to map field theory in curved backgrounds to spin systems in $(1+1)$ dimensions. Since our framework has been shown to work well, a future direction could be its application to problems that remain unsolved in the continuum limit or are difficult to study experimentally. One possibility of interest is the observation of thermal radiation in black hole spacetimes or the inflationary universe. The generalization of our method to interacting fields would also be an interesting direction for future research. In the non-perturbative regime, theoretical calculations of QFT in curved spacetime become complex. Spin systems could provide an experimental approach to studying such unexplored areas in theoretical physics.
@@ Acknowledgements.
We would like to thank Tadashi Kadowaki, Masahiro Nozaki and Yohichi Suzuki for useful discussions and comments. The work of K. M. was supported in part by JSPS KAKENHI Grant Nos. JP20K03976, JP21H05186 and JP22H01217. The work of D. Y. was supported by JSPS KAKENHI Grant Nos. 21H05185, 23K22442, 23K25830, 24K06890, and JST PRESTO Grant No. JPMJPR2118. The work of R. Y. was supported by JSPS KAKENHI Nos. 19K14616 and 20H01838.
@@ A.1 Canonical quantization of Majorana fermions
where we have defined the conjugate momenta by right derivatives. Then, Hamiltonian density $\mathcal{H}=\pi_{\Psi}\partial_{t}\Psi+\pi_{\Psi^{\dagger}}\partial_{t}\Psi^{\dagger}-\mathcal{L}$ becomes Eq. (14). Since Eq. (77) cannot be solved in terms of $\partial_{t}\chi$ and $\partial_{t}\chi^{\dagger}$, the current system is singular and Eq. (77) should be regarded as constraints. We define constrains $\phi_{1}$ and $\phi_{2}$ as
@@ B.1 Continuum theory
$\begin{split}&\vec{\phi}_{k}=\begin{pmatrix}u_{k}^{(1)}\,z_{-}^{i\epsilon_{k}^{(2)}\Delta\eta/2}F(\alpha_{-},\beta_{-},\gamma_{-},z_{+})\\ v_{k}^{(1)}\,z_{-}^{-i\epsilon_{k}^{(2)}\Delta\eta/2}F(\alpha_{+}^{\ast},\beta_{+}^{\ast},\gamma_{+}^{\ast},z_{+})\end{pmatrix}z_{+}^{-i\epsilon_{k}^{(1)}\Delta\eta/2}\ ,\\ &\sigma^{x}\vec{\phi}_{-k}^{\ast}=\begin{pmatrix}v_{k}^{(1)}\,z_{-}^{i\epsilon_{k}^{(2)}\Delta\eta/2}F(\alpha_{+},\beta_{+},\gamma_{+},z_{+})\\ u_{k}^{(1)}\,z_{-}^{-i\epsilon_{k}^{(2)}\Delta\eta/2}F(\alpha_{-}^{\ast},\beta_{-}^{\ast},\gamma_{-}^{\ast},z_{+})\end{pmatrix}z_{+}^{i\epsilon_{k}^{(1)}\Delta\eta/2}.\end{split}$ (94)
@@
$\vec{\phi}_{k}\simeq\begin{pmatrix}u_{k}^{(1)}\\ v_{k}^{(1)}\end{pmatrix}e^{-i\omega_{1}\eta}\ ,\quad\sigma^{x}\vec{\phi}_{-k}^{\ast}\simeq\begin{pmatrix}v_{k}^{(1)}\\ u_{k}^{(1)}\end{pmatrix}e^{i\omega_{1}\eta}\ .$ (97)
@@
$\begin{split}&\vec{\phi}_{k}\simeq\begin{pmatrix}u_{k}^{(1)}(\frac{\Gamma(\gamma_{-})\Gamma(\gamma_{-}-\alpha_{-}-\beta_{-})}{\Gamma(\gamma_{-}-\alpha_{-})\Gamma(\gamma_{-}-\beta_{-})}e^{-i\epsilon_{k}^{(2)}\eta}+\frac{\Gamma(\gamma_{-})\Gamma(\alpha_{-}+\beta_{-}-\gamma_{-})}{\Gamma(\alpha_{-})\Gamma(\beta_{-})}e^{i\epsilon_{k}^{(2)}\eta})\\ v_{k}^{(1)}(\frac{\Gamma(\gamma_{+}^{\ast})\Gamma(\gamma_{+}^{\ast}-\alpha_{+}^{\ast}-\beta_{+}^{\ast})}{\Gamma(\gamma_{+}^{\ast}-\alpha_{+}^{\ast})\Gamma(\gamma_{+}^{\ast}-\beta_{+}^{\ast})}e^{i\epsilon_{k}^{(2)}\eta}+\frac{\Gamma(\gamma_{+}^{\ast})\Gamma(\alpha_{+}^{\ast}+\beta_{+}^{\ast}-\gamma_{+}^{\ast})}{\Gamma(\alpha_{+}^{\ast})\Gamma(\beta_{+}^{\ast})}e^{-i\epsilon_{k}^{(2)}\eta})\end{pmatrix}\ ,\\ &\sigma^{x}\vec{\phi}_{-k}^{\ast}\simeq\begin{pmatrix}v_{k}^{(1)}(\frac{\Gamma(\gamma_{+})\Gamma(\gamma_{+}-\alpha_{+}-\beta_{+})}{\Gamma(\gamma_{+}-\alpha_{+})\Gamma(\gamma_{+}-\beta_{+})}e^{-i\epsilon_{k}^{(2)}\eta}+\frac{\Gamma(\gamma_{+})\Gamma(\alpha_{+}+\beta_{+}-\gamma_{+})}{\Gamma(\alpha_{+})\Gamma(\beta_{+})}e^{i\epsilon_{k}^{(2)}\eta})\\ u_{k}^{(1)}(\frac{\Gamma(\gamma_{-}^{\ast})\Gamma(\gamma_{-}^{\ast}-\alpha_{-}^{\ast}-\beta_{-}^{\ast})}{\Gamma(\gamma_{-}^{\ast}-\alpha_{-}^{\ast})\Gamma(\gamma_{-}^{\ast}-\beta_{-}^{\ast})}e^{i\epsilon_{k}^{(2)}\eta}+\frac{\Gamma(\gamma_{-}^{\ast})\Gamma(\alpha_{-}^{\ast}+\beta_{-}^{\ast}-\gamma_{-}^{\ast})}{\Gamma(\alpha_{-}^{\ast})\Gamma(\beta_{-}^{\ast})}e^{-i\epsilon_{k}^{(2)}\eta})\end{pmatrix}\ .\end{split}$ (98)
On the other hand, natural mode functions at $\eta\to\infty$ are given by
@@ B.2 Discrete theory
$z_{k}=\frac{1}{\varepsilon}(1-\cos\kappa-\varepsilon m)\ ,\quad y(k)=\frac{1}{\varepsilon}\sin k\ .$ (103)
@@ Appendix C On the experimental measurements of the number of produced particles in the expanding Universe
$\begin{split}&\langle c_{j}^{\dagger}c_{l}\rangle=\langle\sigma^{-}_{j}\sigma^{+}_{l}\prod_{r\in(j,l)}\sigma^{z}_{r}\rangle,~{}\langle c_{j}c_{l}^{\dagger}\rangle=-\langle\sigma^{+}_{j}\sigma^{-}_{l}\prod_{r\in(j,l)}\sigma^{z}_{r}\rangle\ ,\\ &\langle c_{j}c_{l}\rangle=\textrm{sgn}(j-l)\langle\sigma^{+}_{j}\sigma^{+}_{l}\prod_{r\in(j,l)}\sigma^{z}_{r}\rangle\ ,\\ &\langle c_{j}^{\dagger}c_{l}^{\dagger}\rangle=-\textrm{sgn}(j-l)\langle\sigma^{-}_{j}\sigma^{-}_{l}\prod_{r\in(j,l)}\sigma^{z}_{r}\rangle\ ,\end{split}$ (108)
@@
i.e., the set of sites between $j$- and $l$-th sites (which does not include its endpoints). The right-hand sides of Eq. (108) are $(|j-l|+1)$-point functions. The product of $\sigma^{z}$’s in multi-point functions is rewritten as
@@
$\begin{split}&\langle c_{j}^{\dagger}c_{l}\rangle=-\langle\sigma^{-}_{j}\sigma^{+}_{l}\prod_{r\notin(j,l),r\neq j,l}\sigma^{z}_{r}\rangle\ ,\\ &\langle c_{j}c_{l}^{\dagger}\rangle=\langle\sigma^{+}_{j}\sigma^{-}_{l}\prod_{r\notin(j,l),r\neq j,l}\sigma^{z}_{r}\rangle\ ,\\ &\langle c_{j}c_{l}\rangle=\textrm{sgn}(j-l)\langle\sigma^{+}_{j}\sigma^{+}_{l}\prod_{r\notin(j,l),r\neq j,l}\sigma^{z}_{r}\rangle\ ,\\ &\langle c_{j}^{\dagger}c_{l}^{\dagger}\rangle=-\textrm{sgn}(j-l)\langle\sigma^{-}_{j}\sigma^{-}_{l}\prod_{r\notin(j,l),r\neq j,l}\sigma^{z}_{r}\rangle\ ,\end{split}$ (111)

### v2

@@ III.1 Spin models and their continuum limits
where $k$ is a typical wave number 333 The long-wavelength limit of the spin system does not imply the low-energy limit in general. For example, in the XY-model, the gap closes twice in the momentum space and the Dirac fermion appears instead of the Majorana fermion in the low-energy limit. In this paper, we will focus on the long-wavelength limit and consider the Majorana field. . Since the physical system size is given by $\ell$, the typical wave number should satisfy $k\gtrsim 1/\ell$. Similarly, we can express all terms in Eq. (21) using $\Psi(x_{j})$. We require that $J^{ab}_{j}(t)=\mathcal{O}(\varepsilon^{-1})$, $h_{j}(t)=\mathcal{O}(\varepsilon^{-1})$ and $J^{+-}_{j}(t)+J^{-+}_{j}(t)-2h_{j}(t)=\mathcal{O}(\varepsilon^{0})$. In terms of their limiting behaviors, we define the real functions $v(t,x)$, $w(t,x)$, $p(t,x)$, $q(t,x)$ and $r(t,x)$ as follows:
@@ III.2 On the choice of free functions: $p(t,x)$ and $\zeta(t,x)$
Note that for a specific choice of $p(t,x)$, such as $p=0$ in the case of the Minkowski metric with $\alpha=\gamma=1$ and $\beta=0$, additional fermionic modes can appear at momenta away from $k\sim 0$ in the low-energy regime, reminiscent of the fermion doubling problem identified by Kenneth Wilson. This implies that the function $p(t,x)$ can be used to control and suppress such doubling, effectively playing the same role as Wilson’s term in the lattice gauge theory [33, 34, 35]. Remarkably, in our dictionary, this term is not inserted by hand but emerges naturally from the structure of the mapping itself.
@@ IV.3 Exactly solvable example
This satisfies $a(\eta)\to a_{1}$ ($\eta\to-\infty$) and $a(\eta)\to a_{2}$ ($\eta\to\infty$). In this case, we have analytical solutions of mode functions both in continuum and discrete theories. The detailed calculation is summarized in appendix B. Once we have two sets of mode functions, we can compute the number of produced particles for a given wave number using Eqs. (54) and (65). As the quantum state $|\Omega\rangle$, we take the ground state at the sufficiently early time. Figure 1 shows the number of particles produced by the cosmic expansion with $m=a_{1}=1$, $a_{2}=2$ and $\ell/2\pi=5$. The upper and lower panels are for $\Delta\eta=1$ and $\Delta\eta=0.2$. Both results for the continuum and discrete theories are shown together. For the discrete theory, we take $k=\kappa/\varepsilon$ as the horizontal axis for the visibility of the continuum limit $L\to\infty$. The number of sites is varied as $L=64,128,256,512$. Since we consider that $\ell$ is fixed, the lattice spacing $\varepsilon=\ell/L$ depends on $L$ while the wave numbers $k$ are defined at the same discrete points for any $L$. In Fig. 1, we plot $n_{k}$ as a function of the wave number at the discrete points represented by dots, together with a continuum variable for visibility. We find that the spectrum tends to converge to that of the continuum theory as $L$ increases. Indeed, the difference between the number of particles produced in discrete and continuum theories, $|n_{k}(L)-n_{k}^{\text{QFT}}|$, depends on $O(1/L)$, as shown in Fig. 2. Especially, for $L\gtrsim 100$, the spectrum qualitatively agrees with that of the continuum theory. This result demonstrates that the spin model can work as a simulator of QFT in the curved spacetime. As the timescale of the cosmic expansion $\Delta\eta$ decreases, more particles are produced. Also, typical wave number of produced particles increases as $\Delta\eta$ decreases. Now, a physical length scale of the systems is given by $\ell$. A typical Hubble scale is $H_{\eta}\equiv a^{-2}da/d\eta\approx 1/a\Delta\eta$ during the cosmic expansion. Therefore, the present phenomena should be governed by the relation between the values of two dimensionless parameters, the mass scale $m\ell$ and the Hubble scale $H_{\eta}\ell$.
@@ V.3 Confirmation of the Unruh effect through the spin system
where we have used $I(r^{\prime})=(\pi/2)(1+r^{\prime}{}^{2}/4+\cdots)=(\pi/2)(1+\varepsilon m/2+\cdots)$. These are the metric functions of the Rindler spacetime with $\kappa=2\pi T_{U}$, as shown in (71). Thus, the entanglement Hamiltonian (69) is directly related to the Rindler Hamiltonian $H_{R}$ in the continuum limit and exhibits the Unruh effect. Note that a related discussion on the XXZ chain has been presented in [43].
@@ VI Summary and discussion
In this work, we established a complete dictionary to map field theory in curved backgrounds to spin systems in $(1+1)$ dimensions. Since our framework has been shown to work well, a future direction could be its application to problems that remain unsolved in the continuum limit or are difficult to study experimentally. One possibility of interest is the observation of thermal radiation in black hole spacetimes or the inflationary universe. The generalization of our method to interacting fields would also be an interesting direction for future research. In the non-perturbative regime, theoretical calculations of QFT in curved spacetime become complex. Spin systems could provide an experimental approach to studying such unexplored areas in theoretical physics. (For example, the dynamics of self-interacting fermionic fields in an expanding universe has recently been discussed with ultracold atoms in optical lattices [49].)
@@ Acknowledgements.
We would like to thank Tadashi Kadowaki, Masahiro Nozaki, Yohichi Suzuki, and Etsuko Itou for useful discussions and comments. The work of K. M. was supported in part by JSPS KAKENHI Grant Nos. JP20K03976, JP21H05186 and JP22H01217. The work of D. Y. was supported by JSPS KAKENHI Grant Nos. 21H05185, 23K22442, 23K25830, 24K06890, and JST PRESTO Grant Nos. JPMJPR2118 and JPMJPR245D. The work of R. Y. was supported by JSPS KAKENHI Nos. 19K14616, 20H01838, and 25K07156.
@@ A.1 Canonical quantization of Majorana fermions
where we have defined the conjugate momenta by right derivatives. Then, Hamiltonian density $\mathcal{H}=\pi_{\Psi}\partial_{t}\Psi+\pi_{\Psi^{\dagger}}\partial_{t}\Psi^{\dagger}-\mathcal{L}$ becomes Eq. (14). Since Eq. (77) cannot be solved in terms of $\partial_{t}\Psi$ and $\partial_{t}\Psi^{\dagger}$, the current system is singular and Eq. (77) should be regarded as constraints. We define constrains $\phi_{1}$ and $\phi_{2}$ as
@@ B.1 Continuum theory
$\begin{split}&\vec{\phi}_{k}=\begin{pmatrix}u_{k}^{(1)}\,z_{-}^{i\epsilon_{k}^{(2)}\Delta\eta/2}F(\alpha_{-},\beta_{-},\gamma_{-},z_{+})\\ v_{k}^{(1)}\,z_{-}^{-i\epsilon_{k}^{(2)}\Delta\eta/2}F(\alpha_{+}^{\ast},\beta_{+}^{\ast},\gamma_{+}^{\ast},z_{+})\end{pmatrix}z_{+}^{-i\epsilon_{k}^{(1)}\Delta\eta/2}\ ,\\ &\sigma^{x}\vec{\phi}_{-k}^{\ast}=\begin{pmatrix}v_{k}^{(1)}\,z_{-}^{i\epsilon_{k}^{(2)}\Delta\eta/2}F(\alpha_{+},\beta_{+},\gamma_{+},z_{+})\\ u_{k}^{(1)}\,z_{-}^{-i\epsilon_{k}^{(2)}\Delta\eta/2}F(\alpha_{-}^{\ast},\beta_{-}^{\ast},\gamma_{-}^{\ast},z_{+})\end{pmatrix}z_{+}^{i\epsilon_{k}^{(1)}\Delta\eta/2},\end{split}$ (94)
@@
$\vec{\phi}_{k}\simeq\begin{pmatrix}u_{k}^{(1)}\\ v_{k}^{(1)}\end{pmatrix}e^{-i\epsilon_{k}^{(1)}\eta}\ ,\quad\sigma^{x}\vec{\phi}_{-k}^{\ast}\simeq\begin{pmatrix}v_{k}^{(1)}\\ u_{k}^{(1)}\end{pmatrix}e^{i\epsilon_{k}^{(1)}\eta}\ .$ (97)
@@
$\begin{split}&\vec{\phi}_{k}\simeq\begin{pmatrix}u_{k}^{(1)}(\frac{\Gamma(\gamma_{-})\Gamma(\gamma_{-}-\alpha_{-}-\beta_{-})}{\Gamma(\gamma_{-}-\alpha_{-})\Gamma(\gamma_{-}-\beta_{-})}e^{-i\epsilon_{k}^{(2)}\eta}+\frac{\Gamma(\gamma_{-})\Gamma(\alpha_{-}+\beta_{-}-\gamma_{-})}{\Gamma(\alpha_{-})\Gamma(\beta_{-})}e^{i\epsilon_{k}^{(2)}\eta})\\ v_{k}^{(1)}(\frac{\Gamma(\gamma_{+}^{\ast})\Gamma(\gamma_{+}^{\ast}-\alpha_{+}^{\ast}-\beta_{+}^{\ast})}{\Gamma(\gamma_{+}^{\ast}-\alpha_{+}^{\ast})\Gamma(\gamma_{+}^{\ast}-\beta_{+}^{\ast})}e^{i\epsilon_{k}^{(2)}\eta}+\frac{\Gamma(\gamma_{+}^{\ast})\Gamma(\alpha_{+}^{\ast}+\beta_{+}^{\ast}-\gamma_{+}^{\ast})}{\Gamma(\alpha_{+}^{\ast})\Gamma(\beta_{+}^{\ast})}e^{-i\epsilon_{k}^{(2)}\eta})\end{pmatrix}\ ,\\ &\sigma^{x}\vec{\phi}_{-k}^{\ast}\simeq\begin{pmatrix}v_{k}^{(1)}(\frac{\Gamma(\gamma_{+})\Gamma(\gamma_{+}-\alpha_{+}-\beta_{+})}{\Gamma(\gamma_{+}-\alpha_{+})\Gamma(\gamma_{+}-\beta_{+})}e^{-i\epsilon_{k}^{(2)}\eta}+\frac{\Gamma(\gamma_{+})\Gamma(\alpha_{+}+\beta_{+}-\gamma_{+})}{\Gamma(\alpha_{+})\Gamma(\beta_{+})}e^{i\epsilon_{k}^{(2)}\eta})\\ u_{k}^{(1)}(\frac{\Gamma(\gamma_{-}^{\ast})\Gamma(\gamma_{-}^{\ast}-\alpha_{-}^{\ast}-\beta_{-}^{\ast})}{\Gamma(\gamma_{-}^{\ast}-\alpha_{-}^{\ast})\Gamma(\gamma_{-}^{\ast}-\beta_{-}^{\ast})}e^{i\epsilon_{k}^{(2)}\eta}+\frac{\Gamma(\gamma_{-}^{\ast})\Gamma(\alpha_{-}^{\ast}+\beta_{-}^{\ast}-\gamma_{-}^{\ast})}{\Gamma(\alpha_{-}^{\ast})\Gamma(\beta_{-}^{\ast})}e^{-i\epsilon_{k}^{(2)}\eta})\end{pmatrix}\ ,\end{split}$ (98)
where $\Gamma(x)$ denotes the Gamma function. On the other hand, natural mode functions at $\eta\to\infty$ are given by
@@ B.2 Discrete theory
$z_{\kappa}=\frac{1}{\varepsilon}(1-\cos\kappa-\varepsilon m)\ ,\quad y_{\kappa}=\frac{1}{\varepsilon}\sin\kappa\ .$ (103)
@@ Appendix C On the experimental measurements of the number of produced particles in the expanding Universe
$\begin{split}&\langle c_{j}^{\dagger}c_{l}\rangle=\langle\sigma^{-}_{j}\sigma^{+}_{l}\prod_{r\in(j,l)}\sigma^{z}_{r}\rangle,\\ &\langle c_{j}c_{l}^{\dagger}\rangle=(-1+2\delta_{jl})\langle\sigma^{+}_{j}\sigma^{-}_{l}\prod_{r\in(j,l)}\sigma^{z}_{r}\rangle\ ,\\ &\langle c_{j}c_{l}\rangle=\textrm{sgn}(j-l)\langle\sigma^{+}_{j}\sigma^{+}_{l}\prod_{r\in(j,l)}\sigma^{z}_{r}\rangle\ ,\\ &\langle c_{j}^{\dagger}c_{l}^{\dagger}\rangle=-\textrm{sgn}(j-l)\langle\sigma^{-}_{j}\sigma^{-}_{l}\prod_{r\in(j,l)}\sigma^{z}_{r}\rangle\ ,\end{split}$ (108)
@@
i.e., the set of sites between $j$- and $l$-th sites (which does not include its endpoints). When $|j-l|\leq 1$, $(j,l)$ is the empty set. When $(j,l)$ is empty, we define $\prod_{r\in(j,l)}\sigma^{z}_{r}=1$. The right-hand sides of Eq. (108) are $(|j-l|+1)$-point functions. The product of $\sigma^{z}$’s in multi-point functions is rewritten as
@@
$\begin{split}&\langle c_{j}^{\dagger}c_{l}\rangle=-\langle\sigma^{-}_{j}\sigma^{+}_{l}\prod_{r\notin(j,l),r\neq j,l}\sigma^{z}_{r}\rangle\ ,\\ &\langle c_{j}c_{l}^{\dagger}\rangle=\langle\sigma^{+}_{j}\sigma^{-}_{l}\prod_{r\notin(j,l),r\neq j,l}\sigma^{z}_{r}\rangle\ ,\\ &\langle c_{j}c_{l}\rangle=\textrm{sgn}(j-l)\langle\sigma^{+}_{j}\sigma^{+}_{l}\prod_{r\notin(j,l),r\neq j,l}\sigma^{z}_{r}\rangle\ ,\\ &\langle c_{j}^{\dagger}c_{l}^{\dagger}\rangle=-\textrm{sgn}(j-l)\langle\sigma^{-}_{j}\sigma^{-}_{l}\prod_{r\notin(j,l),r\neq j,l}\sigma^{z}_{r}\rangle\ .\end{split}$ (111)